The sum of two consecutive even integers is 86. Find the larger number.
Found 2 solutions by JimboP1977, Alan3354:
Answer by JimboP1977(308)

Let x be an even integer
Let x+2 be the next consecutive even integer
x+(x+2) = 86
2x+2 =86
2x = 86-2
Can you solve now for x and find the two even integers?
Post back if you can't.

Answer by Alan3354(21605)

The easy way:
86/2 = 43
One number is just below = 42
One is just above = 44